# Multiple Invertible and Partial-Equivariant Function for Latent Vector Transformation to Enhance Disentanglement in VAEs

### v1

@@ 3.3 Integration for Multiple IPE-Transformation and EF-Conversion
Table 3: p-value of t-test for original vs MIPET results of Table 2, which are averaged over models (bold: positive and significant, italic: positive but insignificant, normal: lower performance).
p-value VAEs CLG-VAE $\beta$-TCVAEs
FVM MIG SAP DCI FVM MIG SAP DCI FVM MIG SAP DCI
dSprites 0.000 0.000 0.000 0.000 0.030 0.000 0.005 0.000 0.281 0.000 0.170 0.009
3D Shapes 0.080 0.007 0.016 0.191 0.085 0.029 0.000 0.088 0.111 0.383 0.277 0.390
3D Cars 0.659 0.250 0.003 0.583 0.003 0.000 0.630 0.071 0.278 0.923 0.119 0.933
4 Experiment Settings
4.0.1 Models
As baseline models, we select VAE, $\beta$-VAE, $\beta$-TCVAE, and CLG-VAE. These models are compared to their extension to adopt MIPET, abbreviated by adding the MIPET prefix. We apply the proposed method to $\beta$-TCVAE only with the EF similarity loss term because $\beta$-TCVAE penalizes the divided KL divergence terms. We set the same encoder and decoder architecture in each model to exclude the overlapped effects. Also, we follow the same model architecture which are introduced in previous works Kim & Mnih (2018) and model details are in Table 8-8.
4.0.2 Datasets
We compare well-known VAEs to CHIC-VAEs: VAE, $\beta$-VAE, $\beta$-TCVAE, and CLG-VAE on the following data sets with 1) dSprites Matthey et al. (2017) which consists of 737,280 binary $64\times 64$ images of dSprites with five independent ground truth factors(number of values), $i.e.$ shape(3), orientation(40), scale(6), x-position(32), and y-position(32). 2) 3D Shapes Burgess & Kim (2018) which consists of 480,000 RGB $64\times 64\times 3$ images of 3D Shapes with six independent ground truth factors: shape(4) orientation(15), scale(8), wall color(10), floor color(10), and object color(10). 3) 3D Cars Reed et al. (2015) which consists of 17,568 RGB $64\times 64\times 3$ images of 3D Shapes with three independent ground truth factors: car models(183), azimuth directions(24), and elevations(4).
4.0.3 Training
We set 256 mini-batch size in the datasets (dSprites, 3D Shapes, and 3D Cars), Adam optimizer with learning rate $4\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and epochs from $\{30,67,200\}$ as a common setting for all the comparative methods. For the comparison, we follow training and inference on the whole dataset. We train each model for 30, 67, and 200 epochs on the dSprites, 3D Shapes, and 3D Cars, respectively, as introduced in Kim & Mnih (2018); Ren et al. (2022). We tune $\beta$ from $\{1,2,4,10\}$ and $\{4,6\}$ for $\beta$-VAE and $\beta$-TCVAE, respectively. We set the dimension size of the latent vectors from $\{6,10\}$ for 10 on dSprites and 3D Cars datasets and 6 for 3D Shapes, but we set 10 for CLG-VAE because it sets 10 dimensions size on 3D Shapes in Zhu et al. (2021). Regarding the CLG-VAE, we fix $\lambda_{\text{decomp}}$, $\lambda_{\text{hessian}}$, and forward group features as 40, 20, and 0.2, respectively. Because the hyper-parameters showed the best result in Zhu et al. (2021). We set group reconstruction from $\{0.2,0.5,0.7\}$. For Control-VAE, we set the maximum KL divergence value from $\{10,12,\ldots,20\}$. In addition, we set masking ratio $\lambda$ from $\{0.0,0.5,\ldots 2.0,\infty\}$. To check the impact of MIPE-transformation, we do not consider the Groupified VAE because the latter is implemented with an extended decoder (different capacity).
4.0.4 Evaluation
We conduct experiments on NVIDIA A100, RTX 2080 Ti, and RTX 3090. We set 100 samples to evaluate global empirical variance in each dimension and run it a total of 800 times to estimate the FVM score introduced in Kim & Mnih (2018). For the MIG Chen et al. (2018), SAP Kumar et al. (2018), and DCI Eastwood & Williams (2018), we follow default values introduced in Michlo (2021), training and evaluation 100 and 50 times with 100 mini-batches, respectively. We evaluate four disentanglement metrics for a less biased understanding of the actual states of disentanglement.
dataset Metrics
FVM IMG SAP DCI
dSprites [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 5: Impact of the mask (mean±$\pm$±std.) and its ratio $\lambda$ in Eq. 8 on 3D Cars. ($\infty$: no masking case, gray box: the best setting over all metrics, bold text: the best in each metric.) Each model runs with ten random seeds.
@@ 4.0.4 Evaluation
$\lambda$ FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$ FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DIC $\uparrow$
@@ 5.1 Quantitative Analysis
5.1.1 Disentanglement Metrics
We set the number of IPE-transformation functions to be equal to balancing hyper-parameter $\beta$ on $\beta$-VAE because of Eq. 11. The number of IPE-transform functions of $\beta$-TCVAE is 3. However, in the case of CLG-VAE, we set it to 1 because its approach is based on the group theory, not directly controlling a KL divergence term such as $\beta$-VAE. We average each model performance value with 40, 20, 60, and 30 cases in VAEs, $\beta$-TCVAEs, Control-VAE and CLG-VAEs, respectively.
@@ 5.1.1 Disentanglement Metrics
We additionally estimate the p-value of each metrics over models in Table 3. Previous work shows the average case of each models Yang et al. (2022).
[…]
Table 6: Ablation study for the equivariant property (w/o E), and EF-conversion (w/o EF). Each metric is averaged over 40 and 20 settings of $\beta$-VAE and $\beta$-TCVAE, respectively.
3D Shapes $\beta$-VAE $\beta$-TCVAE
MIPET MIPET (w/o E) MIPET (w/o EF) MIPET MIPET (w/o E) MIPET (w/o EF)
FVM 75.19($\pm$8.16) 74.91($\pm$10.46) 22.27($\pm$1.29) 80.59($\pm$8.57) 77.90($\pm$8.66) 66.38($\pm$7.57)
MIG 47.37($\pm$10.13) 47.45($\pm$8.98) 0.28($\pm$0.09) 54.49($\pm$9.44) 51.37($\pm$11.54) 36.08($\pm$17.42)
SAP 9.20($\pm$2.44) 9.43($\pm$2.59) 0.26($\pm$0.07) 11.58($\pm$3.32) 10.23($\pm$3.13) 7.13($\pm$3.09)
DCI 54.95($\pm$8.99) 54.23($\pm$9.05) 0.10($\pm$0.02) 66.22($\pm$7.32) 61.18($\pm$8.87) 56.85($\pm$11.72)
5.1.2 Sensitivity to the Number of IPE-transformation and EF-conversion
We investigate the impact of the MIPE-transformation function. As presented in Table. 4, MIPE-transformation is better than IPE-transformation for disentanglement learning on each dataset. Indeed, MIPET-$\beta$-VAEs results more generally and clearly show the impact of the MIPE-transfomation function. Our derivation in Section C.1 clearly explains MIPE-transformation impact. This result shows the impact of the multiple uses of IPE-transformation and EF-conversion.
5.1.3 Impact of Implicit Semantic Mask
@@
5.1.4 Ablation Study
@@
Figure 4: Qualitative results on dSprites. The left-side grids are input images and their variants by changing activations of each dimension of latent vectors. The first row shows input images. The right-side table shows matching pre-defined factors of the dataset (red: MIPET, blue: no MIPET).
Figure 5: Qualitative analysis result of $\beta$-VAE (4.0) with 3D Cars.
@@ 5.2 Qualitative Analysis
[…]
In Figure 4, $\beta$-TCVAE struggles with y-position and rotation, as shown on the 6th row, and with scale and shape represented on the 7th row. On the contrary, MIPET-$\beta$-TCVAE separates y-position and rotation factor (10th, and 7th rows), also the activated dimensions of MIPET-$\beta$-TCVAE are not overlapped with each factor. Applied our method on $\beta$-TCVAE shows better disentangled representation on dSprites dataset. These results also show that our proposed method improves disentangled representation learning.
The qualitative analysis with 3D Cars dataset, as show in Figure 5, the left side is the $\beta$-TCVAE result, and it struggles with body, and azimuth factors shown in the 7th row. However, MIPET-$\beta$-TCVAE separates azimuth (6th row) and body (1st row). In particular, MIPET-$\beta$-TCVAE learns color factor (3rd row) which does not exist on $\beta$-TCVAE.
@@ 6 Conclusion
In this paper, we address the problem of injecting inductive bias for learning unsupervised disentangled representations. To build the bias in VAE frameworks, we propose MIPE-transformation composed of 1) IPE-transformation for the benefits of invertibility and partial-equivariant for disentanglement, 2) a training loss and module to adapt unrestricted prior and posterior to an approximated exponential family, and 3) integration of multiple units of IPE-transformation function and EF-conversion for more expressive bias. The method is easily equipped on state-of-the-art VAEs for disentanglement learning and shows significant improvement on dSprites, 3D Shapes, and 3D Cars datasets. We expect that our method can be applied to more VAEs, and extended to downstream applications. Our work is limited to holding partial equivariance of I2L transformation, so more direct methods to induce it can be integrated in the future.
@@ Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (2022R1A2C2012054).
Appendix A Appendix
Appendix B Preliminaries
B.1 Group Theory
Binary operation: Binary operation on a set $S$ is a function that $\ast:S\times S\rightarrow S$, where $\times$ is a cartesian product.
  • closure: $g_{a},g_{b}\in G\Rightarrow g_{a}\ast g_{b}\in G$.
Group action: Let $(G,\ast)$ be a group and set $X$, binary operation $\cdot:G\times X\rightarrow X$, such that following properties:
  • Identity: $e\cdot x=x$, where $e\in G,x\in X$.
  • Compatibility: $\forall g_{a},g_{b}\in G,x\in X,(g_{a}\ast g_{b})\cdot x=g_{a}\cdot(g_{b}\cdot x)$.
Equivariant map: Let $G$ be a group and $X_{1},X_{2}$ be two sets with corresponding group action of $G$ in each sets: $T^{X_{1}}_{g},T^{X_{2}}_{g}$, where $g\in G$. Then a function $f:X_{1}\rightarrow X_{2}$ is equivariant if $f(T^{X_{1}}_{g}\cdot X_{1})=T^{X_{2}}_{g}\cdot f(X_{1})$.
Partial Equivariance:Romero & Lohit (2022): Let subset of $G$ be $\Upsilon\subset G$, then $f$ is a partially equivariant map to $G$:
$f(T^{X_{1}}_{\upsilon}\cdot X_{1})=T^{X_{2}}_{\upsilon}\cdot f(X_{1}),~{}\text{where}~{}\forall\upsilon\in\Upsilon.$ (12)
Homomorphsim: Let $(G,\cdot),(H,\circ)$ be two groups. If mapping function $h:G\rightarrow H$, s.t. $h(g_{i}\cdot g_{j})=f(g_{i})\circ f(g_{j})$, then $f$ is called homomorphism.
B.2 Exponential Family
Power density function of the exponential family (PDF) generalized formulation:
$\begin{split}f_{{\bm{x}}}({\bm{x}}|\bm{\theta})&=h({\bm{x}})\text{exp}(\bm{\theta}^{\intercal}T({\bm{x}})-A(\bm{\theta}))\\ &=\text{exp}(\bm{\theta}^{\intercal}T({\bm{x}})-A(\bm{\theta})+B({\bm{x}})),\end{split}$ (13)
Appendix C Proof
Proposition C.1.

### v3

@@ 4.2 Exponential Family Conversion for Unknown Prior
Implicit Semantic Mask
where $\mathcal{W}$ is the weight of log-normalizer, $\lambda$ is a hyper-parameter, $\mathbf{\mu}_{|\mathcal{W}_{ij}|}$, and $\mathbf{\sigma}_{|\mathcal{W}_{ij}|}$ are the mean, and standard deviation of weight respectively. Previous work (Yang et al., 2020) utilizes a semantic mask in input space directly, but we inject the semantic mask implicitly on the latent space.
4.3 Integration for Multiple IPE-Transformation and EF-Conversion
$\begin{aligned} \mathcal{L}_{\text{rec}}&=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}}))\end{aligned}$
where $\hat{{\bm{z}}}_{i}=\psi_{i}({\bm{z}})$. Therefore, we define ELBO as:
$\begin{split}\mathcal{L}^{\prime}(\phi,\theta,\psi_{i\in[1,k]};{\bm{x}})&=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi,\psi_{i}}({\bm{z}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}}))\\ &-\sum_{i=1}^{k}D_{\mathrm{KL}}(q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})||p_{\psi_{i}}({\bm{z}})).\end{split}$ (15)
However, following Eq. 15, k samples are generated, and each sample is disentangled for different factors. We implement the output as the average of the sum of the k samples to obtain a single sample with a superposition effect from k samples. Moreover, the KL divergence term in Eq. 15 represents that increasing number of MIPE-transformation is equal to an increasing $\beta$ hyper-parameter in $\beta$-VAE (Higgins et al., 2017) and more details are in Appendix C.
Table 2: Impact of the number of MIPE-transformation function on the $\beta$-TCVAE and $\beta$-VAE with dSprites, 3D Shapes, and 3D Cars datasets in terms of the four metrics. The blue and red box plots represent each model’s single and multiple IPE-transformation cases, respectively (A-$n$: MIPET-$\beta$-TCVAE (4), B-$n$: MIPET-$\beta$-TCVAE (6), C-$n$: MIPET-$\beta$-VAE, $n$: the number of MIPE-transformation). Further details are in Table 9.
dSprites 3D Shapes
FVM MIG SAP DCI FVM MIG SAP DCI
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[…]
@@ 4.3 Integration for Multiple IPE-Transformation and EF-Conversion
$\lambda$ FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$ FVM $\uparrow$ MIG $\uparrow$ SAP $\uparrow$ DCI $\uparrow$
@@
5 Experiment Settings
5.1 Models
As baseline models, we select VAE (Kingma and Welling, 2013), $\beta$-VAE (Higgins et al., 2017), $\beta$-TCVAE (Chen et al., 2018), Factor-VAE (Kim and Mnih, 2018), Control-VAE (Shao et al., 2020),and CLG-VAE (Zhu et al., 2021). These models are compared to their extension to adopt MIPET, abbreviated by adding the MIPET prefix. Also, we follow the same model architecture which are introduced in previous works (Kim and Mnih, 2018) and model details are in Table 7-7.
5.2 Datasets
As in prior works (Higgins et al., 2017; Chen et al., 2018; Kim and Mnih, 2018; Zhu et al., 2021; Yang et al., 2022; Keller and Welling, 2021), we primarily evaluate our models on synthetic datasets, where learning disentangled representations remains a well-known challenge (Matthey et al., 2017; Reed et al., 2015; Burgess and Kim, 2018). We benchmark MIPET-VAEs against several standard VAEs on the dSprites (Matthey et al., 2017), 3D Shapes (Burgess and Kim, 2018), and 3D Cars (Reed et al., 2015) datasets. Further details are in the Appendix D.2
5.3 Training
We set 256 mini-batch size in the datasets (dSprites, 3D Shapes, and 3D Cars), Adam optimizer with learning rate $4\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and epochs from $\{30,67,200\}$ as a common setting for all the comparative methods. In addition, we set masking ratio $\lambda$ from $\{0.0,0.5,\ldots 2.0,\infty\}$. We then evaluate familiar benchmarks, FVM score (Kim and Mnih, 2018), MIG (Chen et al., 2018), SAP (Kumar et al., 2018), and DCI (Eastwood and Williams, 2018). Further details are in the Appendix D.3.
Table 4: Ablation study for the equivariant property (w/o E) and EF-conversion (w/o EF). Each metric is averaged over 40 and 20 settings of $\beta$-VAE and $\beta$-TCVAE, respectively.
3D Shapes $\beta$-VAE $\beta$-TCVAE Factor-VAE CLG-VAE
MIPET w/o E w/o EF MIPET w/o E w/o EF MIPET w/o E w/o EF MIPET w/o E w/o EF
FVM 75.19($\pm$8.16) 74.91($\pm$10.46) 22.27($\pm$1.29) 80.59($\pm$8.57) 77.90($\pm$8.66) 66.38($\pm$7.57) 78.34($\pm$8.35) 71.50($\pm$6.78) 59.28($\pm$5.15) 80.17($\pm$8.43) 69.75($\pm$6.91) 61.55($\pm$9.05)
MIG 47.37($\pm$10.13) 47.45($\pm$8.98) 0.28($\pm$0.09) 54.49($\pm$9.44) 51.37($\pm$11.54) 36.08($\pm$17.42) 48.59($\pm$8.35) 42.25($\pm$11.28) 32.67($\pm$10.49) 53.87($\pm$7.41) 45.92($\pm$4.77) 33.89($\pm$3.46)
SAP 9.20($\pm$2.44) 9.43($\pm$2.59) 0.26($\pm$0.07) 11.58($\pm$3.32) 10.23($\pm$3.13) 7.13($\pm$3.09) 9.84($\pm$2.73) 8.19($\pm$0.69) 7.04($\pm$0.87) 12.83($\pm$3.01) 50.14($\pm$5.43) 46.66($\pm$4.81)
DCI 54.95($\pm$8.99) 54.23($\pm$9.05) 0.10($\pm$0.02) 66.22($\pm$7.32) 61.18($\pm$8.87) 56.85($\pm$11.72) 60.23($\pm$9.90) 50.65($\pm$6.85) 48.83($\pm$5.37) 60.74($\pm$7.77) 50.14($\pm$5.43) 46.66($\pm$4.81)
Figure 3: Each square represents a value in the DCI matrix, which describes the relationship between the $i^{th}$ latent dimension and each factor. The size of each square is relative to the values within each row. The ideal case resembles a sparse matrix. The y-axis represents the factors of each dataset, while the x-axis corresponds to the latent vector dimensions. The number shown in each row of the matrix indicates the maximum value and standard deviation of that row. Higher maximum and standard deviation values suggest greater sparsity, indicating closer alignment with the ideal case.
(a) Posterior visualization with Beta distribution
(b) Posterior visualization of MIPET layer
Figure 4: Non-Gaussian posterior learned by an IPE module without intended guidance to a specific distribution in a toy setting. We compare VAE, MIPET-VAE, and MIPET-VAE without a semantic mask to assess how well each model captures the underlying distribution. We construct the VAE with a 4-layer Multi-Layer Perceptron (MLP) as the encoder and a single linear layer as the decoder. Blue plots are randomly sampled from a two-dimensional beta distribution, red plots are the posterior, and black plots are the output results.
@@ 6.1 Quantitative Analysis
Disentanglement Metrics
We set the number of IPE-transformation functions to be equal to balancing hyper-parameter $\beta$ on $\beta$-VAE because of Eq. 16. The number of IPE-transform functions of $\beta$-TCVAE is 3. However, in the case of CLG-VAE, we set it to 1 because its approach is based on the group theory, not directly controlling a KL divergence term such as $\beta$-VAE. We average each model performance value with 40, 20, 60, 10, and 30 cases in VAEs, $\beta$-TCVAEs, Control-VAE, Factor-VAE and CLG-VAEs, respectively.
@@
Sensitivity to the Number of IPE-transformation and EF-conversion
We analyze the impact of the MIPE-transformation function. As presented in Table 2, MIPE-transformation outperforms IPE-transformation in disentanglement learning across all datasets. Indeed, the results of MIPET-$\beta$-VAEs more generally and clearly demonstrate the impact of the MIPE-transfomation function. Our derivation in Section C clearly explains the impact of MIPE-transformation. This result shows the impact of the multiple uses of IPE-transformation and EF-conversion.
Impact of Implicit Semantic Mask
@@
Figure 5: Qualitative results on various datasets, which show the factors learned for each dimension of $z$.
Ablation Study
@@ 6.2 Qualitative Analysis
Relation Between Factors and Latent Vector Dimensions
As shown in Fig. 3, our model shows a close to sparse matrix compared to the $\beta$-VAE result in 3D Shapes, also the maximum value and standard deviation of each row are larger than $\beta$-VAE. This indicates that our model achieves better alignment of individual factors with specific latent dimensions compared to other methods. More details are in Appendix D.4.
The Impact of IPE-transformation
The IPE-transformation $\psi(\cdot)$ is designed to achieve two objectives: 1) to represent a non-Gaussian exponential family distribution as the posterior, and 2) to preserve the inherent structure of the dataset. To assess the first objective, we construct a toy dataset sampled from a beta distribution (additional cases are provided in Appendix D.7) and configure the decoder as a single linear layer, noting that the linear transformation of a Beta distribution remains within the same family. As shown in Fig. 4(a), the posterior learned by MIPET-VAE (red dots) closely approximates the beta distribution, whereas the posteriors of VAE and MIPET (without the semantic mask) retain Gaussian distribution. Regarding the second objective, MIPET-VAE exhibits improved preservation of the dataset structure relative to VAE (black dots). Additional experimental details are reported in Appendix D.7.
Does Each IPE-transformation Independently Operate?
We demonstrate how each IPE-transformation represents the exponential family. As shown in Fig. 4(b), each IPE-transformation represents a different distribution because the green and red plots are distributed in different posteriors, even though all IPE-transformations share the same decoder weights.
Disentanglement of Factors Across Latent Dimensions
We randomly sample an image for each dimension of the latent vector space and creates 5 variants of its generated latent vector by selecting values from $\{$-2, 2}\} with 5 intervals for the dimension, then generate their corresponding output images. For the generation, we select $\beta$-TCVAE (6), which shows the best FVM scores in dSprites dataset. Thereafter, we evaluate the semantic roles of each dimension before and after applying MIPE-transformation function.
In Fig. 5(a), $\beta$-TCVAE struggles with rotation and y-position, as shown on the 3rd row, and with scale and shape represented on the 4th row. On the contrary, MIPET-$\beta$-TCVAE separates rotation and y-position factor (4th, and 5th rows). Applied our method on $\beta$-TCVAE shows better disentangled representation on dSprites dataset. These results also show that our proposed method improves disentangled representation learning. As shown in the Fig. 5(b), CLG-VAE struggles with shape and wall color factors in 2nd dimension, and shape and object color factors in 3rd dimension. In particular, it struggles with three factors in 5th dimension. On the other hand, MIPET-CLG-VAE separates shape, wall, and object color factors. The qualitative analysis with 3D Cars dataset, as shown in Fig. 5(c), the left side is the $\beta$-TCVAE result, and it struggles with azimuth and body factors shown in the 2nd row. However, MIPET-$\beta$-TCVAE separates azimuth (4th row) and body (1st row). In particular, MIPET-$\beta$-TCVAE learns color factor (2rd row) which does not exist on $\beta$-TCVAE. Furthermore, our model shows improvement on the CelebA Liu et al. (2015) dataset, as shown in Fig. 5(d). $\beta$-VAE struggles with most factors of the dataset, but ours separates all factors.
@@ 7 Conclusion
In this paper, we address the problem of injecting inductive bias for learning unsupervised disentangled representations. To build the bias in VAE frameworks, we propose MIPE-transformation composed of 1) IPE-transformation for the benefits of invertibility and partial-equivariance for disentanglement, 2) a training loss and module to adapt unrestricted prior and posterior to an approximated exponential family, and 3) integration of multiple units of IPE-transformation function and EF-conversion for more expressive bias. The method is easily equipped on state-of-the-art VAEs for disentanglement learning and shows significant improvement on dSprites, 3D Shapes, and 3D Cars datasets. We expect that our method can be applied to more VAEs, and extended to downstream applications.
@@ Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.2022R1A2C2012054, Development of AI for Canonicalized Expression of Trained Hypotheses by Resolving Ambiguity in Various Relation Levels of Representation Learning).
Checklist
The checklist follows the references. For each question, choose your answer from the three possible options: Yes, No, Not Applicable. You are encouraged to include a justification to your answer, either by referencing the appropriate section of your paper or providing a brief inline description (1-2 sentences). Please do not modify the questions. Note that the Checklist section does not count towards the page limit. Not including the checklist in the first submission won’t result in desk rejection, although in such case we will ask you to upload it during the author response period and include it in camera ready (if accepted).
In your paper, please delete this instructions block and only keep the Checklist section heading above along with the questions/answers below.
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes - We present in Section 4, and Appendix B.]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes - We present in the Appendix D.6.]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes - We submit our code as supplement materials.]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes - We present in Section 4, and Appendix B.]
    2. (b)
      Complete proofs of all theoretical results. [Yes - We present in Section 4, and Appendix B.]
    3. (c)
      Clear explanations of any assumptions. [Yes - We present in Section 4, and Appendix B.]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes - We submit our code as supplement materials.]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes - We present in Section 5 and Appendix D.]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes - We present in Section 6.]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes - We present in Appendix D.]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes - We submit our code as supplement materials.]
    2. (b)
      The license information of the assets, if applicable. [Yes - We submit our code as supplement materials.]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes - We submit our code as supplement materials.]
    4. (d)
      Information about consent from data providers/curators. [Yes - We present in Appendix D.]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
MIPET Supplementary Materials
Appendix A Notation and Groups
Table 5: Terms and Notations
$\psi(\cdot)$ Invertible function
$\bm{z}$} Latent vector from encoder
$\hat{{\bm{z}}}_{m}$ Transformed latent vector by $\psi_{m}(\cdot)$
$\bm{\theta}_{\hat{{\bm{z}}}_{m}}$ Natural Parameter of posterior
$\hat{\bm{\epsilon}}_{m}$ Transformed prior samples by $\psi_{m}(\cdot)$
$\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}$ Natural Parameter of prior
AA Log-Normalizer
$T$ Sufficient Statistics
$\bm{\nu}$ Evidence
$f_{{\bm{x}}}(\cdot)$ Power Density Function
$GL_{n}(\mathbb{R})$ General Linear Group
$M_{n}(\mathbb{R})$ A set of $n\times n$ real matrix
$Sym_{n}(\mathbb{R})$ A set of $n\times n$ symmetric real matrix
$E_{M}$ $\{\mathbf{e}^{M}|M\in M_{n}(\mathbb{R})\}$
$E_{S}$ $\{\mathbf{e}^{S}|S\in Sym_{n}(\mathbb{R})\}$
$\psi_{M}(\cdot)$ $\psi_{M}(\cdot)\in M_{n}(\mathbb{R})$
$\psi_{E_{S}}(\cdot)$ $\psi_{E_{S}}(\cdot)\in E_{S}$
$\psi_{E_{M}}(\cdot)$ $\psi_{E_{M}}(\cdot)\in E_{M}$
$G_{S}$ $G_{S}:(\mathbf{e}^{S},\ast)$
$G_{I}$ Group of input space for symmetries
$G_{L}$ Group of latent space for symmetries
$G_{T}$ Group of transformed latent space for symmetries
$J$ $G_{S}\cap G_{L}$
$G_{I}^{J}$ Group of input space for symmetries corresponds to $J$
$G_{L}^{J}$ Group of latent space for symmetries corresponds to $J$
$G_{L}^{J}$ Group of transformed latent space for symmetries corresponds to $J$
$\mathcal{X}$ Input space
$\mathcal{Z}$ Latent vector space
$\hat{\mathcal{Z}}$ Transformed latent vector space
$\mathcal{Z}^{\prime}$ Homogeneous space of group $G_{L}$ (homogeneous $G_{L}$-space), $\mathcal{Z}^{\prime}\subseteq\mathcal{Z}$
$\hat{\mathcal{Z}}^{\prime}$ Homogeneous space of group $G_{T}$ (homogeneous $G_{T}$-space), $\hat{\mathcal{Z}}^{\prime}\subseteq\hat{\mathcal{Z}}$
$\hat{\mathcal{Z}}^{J}$ Homogeneous space of group $G_{T}^{J}$ (homogeneous $G_{T}^{J}$-space), $\hat{\mathcal{Z}}^{J}\subseteq\hat{\mathcal{Z}}$
$D_{\mathrm{KL}}(\cdot||\cdot)$ Kullback-Leibler divergence
$\mathbf{0}$ zero vector
$\mathbf{0}_{n,n}$ n by n zero matrix
Notation.
We denote $g_{I}\in G_{I}$, $g_{L}\in G_{L}$, and $g_{T}\in G_{T}$. Let $M_{n}(\mathbb{R})$ be the set of $n\times n$ real matrices and $\mathrm{Sym}_{n}(\mathbb{R})$ the set of real symmetric matrices. For any set $\mathcal{A}\subseteq\mathrm{Sym}_{n}(\mathbb{R})$, define $\exp(\mathcal{A}):=\{\exp(S)\mid S\in\mathcal{A}\}$. We write $E_{M}:=\{\exp(M)\mid M\in M_{n}(\mathbb{R})\}$ and $E_{S}:=\exp(\mathrm{Sym}_{n}(\mathbb{R}))$.
Standing assumption on $G_{S}$. We assume a pairwise-commuting symmetric set $\mathcal{S}\subseteq\mathrm{Sym}_{n}(\mathbb{R})$, i.e., $S_{1},S_{2}]=0$ for all $S_{1},S_{2}\in\mathcal{S}$. Then we define
$G_{S}:=\exp(\mathcal{S})=\{\exp(S)\mid S\in\mathcal{S}\}\subset\mathrm{GL}_{n}(\mathbb{R}),$
with the group operation given by the usual matrix multiplication $\ast$”.
Appendix B Proof
Lemma B.1 (Abelian Lie subgroup).
If $\mathcal{S}\subseteq\mathrm{Sym}_{n}(\mathbb{R})$ is pairwise-commuting, then $(G_{S},\ast)$ is an abelian Lie subgroup of $\mathrm{GL}_{n}(\mathbb{R})$. In particular, $G_{S}$ is closed under multiplication and inversion, has identity $I=\exp(0)$, and is commutative.
Proof sketch.
Since all $S\in\mathcal{S}$ commute and are symmetric, they are simultaneously diagonalizable by some orthogonal matrix $Q$: $S_{i}=Q\Lambda_{i}Q^{\top}$ with diagonal $\Lambda_{i}$. Hence $\exp(S_{i})=Q\exp(\Lambda_{i})Q^{\top}$ are symmetric positive definite and commute. Moreover, for any $S_{1},S_{2}\in\mathcal{S}$, the Baker–Campbell–Hausdorff formula reduces to $\exp(S_{1})\exp(S_{2})=\exp(S_{1}+S_{2})$ because $S_{1},S_{2}]=0$. Since $S_{1}+S_{2}\in\mathcal{S}$, we have closure. The inverse is $\exp(S)^{-1}=\exp(-S)$ and $S\in\mathcal{S}$. Commutativity follows from $S_{1},S_{2}]=0$. Smoothness of $\exp$ implies $G_{S}$ is a (abelian) Lie subgroup. ∎
Proposition B.2.